# Implicit neural representations for accurate estimation of the standard model of white matter

### v1

@@ 2.5 Comparisons
The performance of the INR model (referred to as the INR method) is compared to three other SM model fitting methods described below. A supervised machine learning approach using the SMI toolbox (Coelho et al., 2022) (SMI method) with standard settings. The SMI method requires a noise map, which was determined using MP-PCA denoising (Veraart et al., 2016). A supervised deep learning method (referred to as the supervised NN method introduced in (de Almeida Martins et al., 2021)) was trained using a combination of synthetic and data-driven parameter samples. Specifically, the training data consisted of 500,000 samples, with 75% allocated for training and 25% for validation. Half of the training samples were generated by uniformly sampling model parameters, while the other half were derived by applying mutations to nonlinear least squares (NLLS) estimates obtained from the target dMRI data. The neural network architecture comprised three hidden layers with 150, 80, and 55 neurons. Gaussian noise was added with signal to noise ratio (SNR) 50. For a comprehensive description of all training settings, see (de Almeida Martins et al., 2021). Finally, a NLLS approach (NLLS method) was implemented with the MATLAB (MathWorks, Natick, MA, USA) optimization toolbox: lsgnonlin with Levenberg-Marquardt algorithm, max 1000 iterations.
@@ 2.6 Generation of simulated ground truth data
In silico experiments were conducted on simulated data obtained from one brain (subject 011) of the MGH Connectome Diffusion Microstructure Dataset (Tian et al., 2022a). The dMRI data were acquired on the 3T Connectome MRI scanner (Magnetom CONNECTOM, Siemens Healthineers). The acquisitions with $\text{b}=[0,50,350,800,1500,\\ 2400,3450,4750,6000]$$\mathrm{s/mm^{2}}$ and $\Delta=19$ms were selected. The lowest 4 b-values were acquired with 32 uniformly distributed diffusion encoding directions, the highest 4 b-values with 64. The dataset also contained 50 b=0s/mm2 volumes. More details about the imaging parameters and processing can be found in (Tian et al., 2022b). The SM was fitted with the SMI toolbox (Coelho et al., 2022) to generate a set of realistic SM kernel parameters for $f_{i}$, $D_{i}$, $D_{e}$, and $D_{p}$, using $l_{max}=4$ and noise bias correction with a sigma map acquired through MP-PCA (Veraart et al., 2016). To enhance the smoothness of the kernel maps, anisotropic diffusion filtering was performed using MATLAB’s imdiffusefilt function with three iterations (N=3) and minimal connectivity. The SH coefficients $p_{lm}$ of the FOD were calculated using Multi-Shell Multi-Tissue Constrained Spherical Deconvolution (MSMT CSD) (Jeurissen et al., 2014) for $l_{max}=[2,4,6,8]$. The simulated signals corresponding to these parameters were calculated from the SM signal equation with a published optimized acquisition protocol (Coelho et al., 2022): $\text{b}=[0,1000,2000,8000,5000,2000]\,\mathrm{s/mm^{2}}$, number of directions $[4,20,40,40,35,15]$, and B-tensor shape $b_{\Delta}=[1,1,1,1,0.8,0]$. Directions were optimized by minimization of the electric potential energy on a hemisphere (Garyfallidis et al., 2014) after which half of the directions were flipped. Finally, Gaussian noise was added. The standard deviation of the noise distribution was determined by the mean of the b=0s/mm2 acquisitions and the SNR (20,50 or $\infty$), resulting in a spatially varying standard deviation. Free water contributions and TE dependence were not considered. Non-white matter voxels are masked out as their influence on the loss value will decrease the performance of parameter estimation on white matter voxels. A white matter mask was generated with the Freesurfer (Fischl, 2012) segmentation which is included in the MGH dataset. Voxels with $D_{e}>D_{p}$ were also masked out as these represent nonphysical behavior.
@@ 2.7 In vivo data acquisition
The study was approved by the Cardiff University School of Psychology Ethics Committee and written informed consent was obtained from the participant in the study. One healthy volunteer was scanned on a 3T, 300 mT/m Connectom scanner (Siemens Healthineers, Erlangen, Germany). Imaging parameters and diffusion acquisition scheme can be found in Table 2. Gradient waveforms for spherical tensor encoding were optimized using the NOW toolbox (Sjölund et al., 2015).
@@ 2.8.2 Experiment 2: Rician noise bias
The effect of Rician noise bias was investigated by introducing noise sampled from a Rician distribution (SNR = [20,50]), rather than Gaussian. Parameter estimation was performed using both a standard MSE loss and a Rician likelihood loss (Parker et al., 2023), and the resulting estimations were compared to the ground truth. The integral in the SM forward model is calculated with the analytical approach.
@@ 2.8.3 Experiment 3: Qualitative comparison on in vivo data
To test the INR on in vivo data, SM parameter estimation was executed on the dataset from section 2.7. Only $l_{max}=2$ was considered. As the acquisition protocol contained negative $b_{\Delta}$ values, the SM forward model was calculated following the numerical integration approach. The results are compared to the estimates from the methods in section 2.5.
@@ 2.8.4 Experiment 4: Estimation of SH order up to $l_{max}=8$
The performance of the proposed method to model higher order FODs is investigated by fitting the INR with SH orders of $l_{max}=\{2,4,6,8\}$, for three different datasets. The noiseless synthetic dataset and the SNR 50 synthetic data provide insight in the accuracy of FOD estimation in noisy data, while the in vivo dataset qualitatively evaluates the capability to estimate higher order FODs on realistic datasets.
@@ 2.8.5 Experiment 5: Effect of gradient non-uniformity correction of SM parameter estimation
The impact of gradient non-uniformities on SM parameter estimation was assessed using in vivo data. For each voxel, the b-value, $b_{\Delta}$, and b-vectors were recalculated to account for scanner-specific gradient deviations following (Bammer et al., 2003). These corrected diffusion parameters were then incorporated into the training loop by providing them as additional inputs to the data loader. This analysis was performed for $l_{max}=2$. The differences between the corrected and uncorrected parameter estimates were subsequently evaluated. Gradient non-uniformity correction was implemented using both MSE and Rician loss functions.
@@ 2.8.6 Experiment 6: Implicit neural representation for spatial interpolation
To obtain more detailed insight into the continuous spatial representation of the dataset provided by an INR, the INR fit on the Gaussian noise, SNR 50 synthetic dataset at $l_{max}=2$ is sampled at 8x the original resolution in every dimension, resulting in 0.375mm isotropic voxels. The parameter map of $p_{2}$ was visualized in the coronal and sagittal plane for the original resolution output of the model and the linear, cubic and INR upsampling.
@@ 3.1 Quantitative comparison on simulated data
Parameter estimation on simulated data without noise and with SNR = 50 can be found in B, showing a less prominent, but still evident improvement in $\rho$ and RMSE of the INR method compared to other estimation methods for all parameters. Fitting without noise shows that the ground truth generating process is not positively biased towards parameter estimates of the INR.
@@ 3.2 Rician noise bias
The correlation plots between the INR with MSE or Rician loss and the ground truth parameters, alongside the parameter maps for both approaches can be seen in Figure 4. For MSE, the Rician bias is visible in the scatterplots and parameter maps as a general over- or underestimation across all voxels. The Rician loss is able to correct the bias, resulting in better correlation with the ground truth. The bias is most significant for $D_{i}$ and $D_{e}$. The bias effect is less prominent when considering SNR = 50, see supplementary information Fig. 12.
@@ 3.4 Estimation of SH order up to $l_{max}=8$
The INR outputs for the different SH order FODs are visualized in Figure 6. Qualitative inspection of the FODs shows plausible FOD shapes and directions throughout all datasets and SH orders. Furthermore, there are no notable differences between the FOD estimates for the clean and noisy synthetic datasets, indicating noise robustness in the estimate. The backgrounds in Figure 6 show no large voxel-wise differences for $f_{i}$ across the different orders. This holds true for all kernel parameter estimates, which is shown in further detail in D.
Figure 6: Visualization of the centrum semi-ovale for different SH orders $l_{max}$ (rows) and datasets (columns), with a parameter map of $f_{i}$ as background. FODs are scaled for visibility.
@@ 3.5 Effect of gradient non-uniformity correction of SM parameter estimation
Brain parameter maps with and without gradient non-uniformity correction on in vivo data are presented in Fig. 7. Difference maps show a significant effect on $D_{i}$, $D_{e}$. Corrected maps show effect at the edges of the brain, mainly in the frontal lobe. This is to be expected as gradient non-uniformity is strongest there. $D_{i}$ and $D_{e}$ show lower values after correction. The influence of the correction is least apparent on $f_{i}$ and $p_{2}$. Training with Rician loss shows the most effect on $p_{2}$ estimates.
Figure 7: Effect of gradient non-uniformity correction on the estimation of SM parameters. Top row shows parameter maps without correction. Bottom rows show the effect of gradient non-uniformity correction with MSE loss and Rician loss. Difference maps are with respect to prediction with MSE and without gradient non-uniformity correction (top row).
@@ 4.1 Implications of the results
In this work, we show how INRs can be used to estimate continuous, noise-robust SM parameter maps of simulated and in vivo datasets and with different SH orders. The quantitative experiments on synthetic data show how the proposed method outperforms existing methods on noisy data. For high SNR levels (see Fig. 11), the supervised NN method achieves performance metrics close to those of the INR approach; however, it can be significantly more time-consuming due to its reliance on NLLS for generating part of the training data and the need to retrain the model for each specific acquisition protocol. At higher noise levels (SNR = 20), the INR method clearly outperforms all other methods (see Fig. 3). On in vivo data the underlying representation shows a more structurally correlated appearance, without large inter-voxel variability. This is further illustrated by upsampling the INR at high resolution. Parameter estimates for FOD up to SH order eight can be provided alongside the other SM parameters without introducing bias in other parameters as shown in D. The unsupervised nature of the method avoids training-set bias prevalent in supervised fitting methods.
@@
The proposed hyperparameters $n_{p}=5000$ and $n_{h}=2048$ provide a robust setting that can provide good representations of dMRI datasets with different sizes, acquisition protocols, and levels of noise. An exploration of these hyperparameters can be found in the supplement of (Hendriks et al., 2025). The hyperparameter $\sigma^{2}$ provides a convenient way of tuning the model to provide stronger or weaker spatial regularization, as detailed in A.
@@
Fitting time is a critical factor when applying dMRI microstructure modeling, and various efforts have been made to speed up the computationally heavy nonlinear optimization and enable large-scale population studies (Daducci et al., 2015a; Harms et al., 2017; Diao and Jelescu, 2023). INRs circumvent this through its inherent unsupervised MLP structure, which allows for efficient, continuous representation of the parameter space without requiring voxel-wise optimization. The INRs are fit on consumer grade hardware in around 6 up to at most 23 minutes for the scenarios tested, much faster than classic NLLS approaches. Using the analytical integration approach decreases training time considerably, possible when excluding negative $b_{\Delta}$ values in acquisition protocols. Once the INR is fit to the dataset, the inference time is negligible. For example, an $lmax=2$ model with $n_{h}=2048$ and $n_{p}=5000$ can perform inference at one million coordinates in 2.7 seconds, including data writing times to and from the GPU. However, since INRs require a model to be fit to every individual dataset, supervised inference remains faster for datasets consisting of many acquisition parameter settings, despite the considerable amount of training time it requires, e.g. for the supervised NN method. Nonetheless, the duration of the fitting process for unsupervised learning in combination with dMRI models have been reduced to seconds when applying transfer learning (Li et al., 2024), which could make on-the-fly fitting of INRs for datasets with a high number of acquisitions possible.
@@
The method’s ability to incorporate gradient non-uniformity correction in the fitting process provides an advantage over typical supervised methods, for which the training set would be impractically large to capture the spatial variability. This correction is essential as even small non-uniformities can affect parameter maps (Bammer et al., 2003; Mesri et al., 2020; Mohammadi et al., 2012; Morez et al., 2021), and the availability of high-performance gradient coils suffering from significant gradient non-uniformities is increasing (Jones et al., 2018). To our knowledge, only method 2 has implemented gradient non-uniformity correction in the SM fitting process (Coelho et al., 2023) using SVD and linear regression methods, which approximates the exact acquisition setting in every voxel.
4.2 Limitations of the approach
The ground truth generated for the synthetic experiments makes use of SMI for generating the underlying parameter maps. This could potentially bias the parameters to be in a range that favors estimation using SMI. We indeed observed that estimation on the noiseless signal showed optimal performance for SMI and NLLS (see Fig. 10a). Nevertheless, we have attempted to reduce biases towards the estimation for creating the ground truth by smoothing the parameter maps and using FODs from a different source (MSMT-CSD). Similarly, we used the diffusion parameter distributions found using NLLS on the target dataset to fit the supervised NN to generate the best-case scenario for the distribution of the training dataset. Commonly the supervised NN would depend on parameter distributions from a predefined fixed dataset, because refitting the network for every dataset is prohibitively time-consuming. Another limitation related to the ground truth is that the MGH dataset used in this study contained only linear tensor encoding (LTE), which may have led to inaccurate parameter estimates (see discussion at the end of this section).

### v2

@@ 2.5 Comparisons
The performance of the INR model (referred to as the INR method) is compared to three other SM model fitting methods described below. A supervised machine learning approach using the SMI toolbox (Coelho et al., 2022) (SMI method) with standard settings. The SMI method requires a noise map, which was determined using MP-PCA denoising (Veraart et al., 2016). A supervised deep learning method (referred to as the supervised NN method introduced in (de Almeida Martins et al., 2021)) was trained using a combination of synthetic and data-driven parameter samples. Specifically, the training data consisted of 500,000 samples, with 75% allocated for training and 25% for validation. Half of the training samples were generated by uniformly sampling model parameters, while the other half were derived by applying mutations to nonlinear least squares (NLLS) estimates obtained from the target dMRI data. The neural network architecture comprised three hidden layers with 150, 80, and 55 neurons. Gaussian noise was added with signal to noise ratio (SNR) 50. For a comprehensive description of all training settings, see (de Almeida Martins et al., 2021). Finally, a NLLS approach (NLLS method) was implemented with the MATLAB (MathWorks, Natick, MA, USA) optimization toolbox: lsgnonlin with Levenberg-Marquardt algorithm, max 1000 iterations. Two initializations were fitted after which the solution with lowest residual norm was chosen.
@@ 2.6 Generation of simulated ground truth data
In silico experiments were conducted on simulated data obtained from one brain (subject 011) of the MGH Connectome Diffusion Microstructure Dataset (Tian et al., 2022a). The dMRI data were acquired on the 3T Connectome MRI scanner (Magnetom CONNECTOM, Siemens Healthineers) at 2mm isotropic resolution. The acquisitions with $\text{b}=[0,50,350,800,1500,\\ 2400,3450,4750,6000]$$\mathrm{s/mm^{2}}$ and $\Delta=19$ms were selected. The lowest 4 b-values were acquired with 32 uniformly distributed diffusion encoding directions, the highest 4 b-values with 64. The dataset also contained 50 b=0s/mm2 volumes. More details about the imaging parameters and processing can be found in (Tian et al., 2022b). The SM was fitted with the SMI toolbox (Coelho et al., 2022) to generate a set of realistic SM kernel parameters for $f_{i}$, $D_{i}$, $D_{e}$, and $D_{p}$, using $l_{max}=4$ and noise bias correction with a sigma map acquired through MP-PCA (Veraart et al., 2016). Further settings were 2 compartments (intra- and extra-axonal), $10^{6}$ training samples and 1 Nlevels. To enhance the smoothness of the kernel maps, anisotropic diffusion filtering was performed using MATLAB’s imdiffusefilt function with three iterations (N=3) and minimal connectivity. The SH coefficients $p_{lm}$ of the FOD were calculated using Multi-Shell Multi-Tissue Constrained Spherical Deconvolution (MSMT CSD) (Jeurissen et al., 2014) for $l_{max}=[2,4,6,8]$. The simulated signals corresponding to these parameters were calculated from the SM signal equation with a published optimized acquisition protocol (Coelho et al., 2022): $\text{b}=[0,1000,2000,8000,5000,2000]\,\mathrm{s/mm^{2}}$, number of directions $[4,20,40,40,35,15]$, and B-tensor shape $b_{\Delta}=[1,1,1,1,0.8,0]$. Directions were optimized by minimization of the electric potential energy on a hemisphere (Garyfallidis et al., 2014) after which half of the directions were flipped. The image resolution was kept identical to the original dataset at 2mm isotropic. Finally, Gaussian or Rician noise was added. The standard deviation of the noise distribution was determined by the mean of the b=0s/mm2 acquisitions and the SNR (20,50 or $\infty$ on the b=0s/mm2 images), resulting in a spatially varying standard deviation. Free water contributions and TE dependence were not considered. Non-white matter voxels are masked out as their influence on the loss value will decrease the performance of parameter estimation on white matter voxels. A white matter mask was generated with the Freesurfer (Fischl, 2012) segmentation which is included in the MGH dataset. Voxels with $D_{e}>D_{p}$ were also masked out as these represent nonphysical behavior.
@@ 2.7 In vivo data acquisition
The study was approved by the Cardiff University School of Psychology Ethics Committee and written informed consent was obtained from the participant in the study. One healthy volunteer was scanned on a 3T, 300 mT/m Connectom scanner (Siemens Healthineers, Erlangen, Germany). Imaging parameters and diffusion acquisition scheme can be found in Table 2. Gradient waveforms for spherical tensor encoding were optimized using the NOW toolbox (Sjölund et al., 2015). The in vivo data was corrected for Gibbs ringing (Kellner et al., 2016), signal drift (Vos et al., 2017), motion and eddy current correction (Nilsson et al., 2015), susceptibility correction (Andersson et al., 2003), and gradient non-uniformity image distortion and B-matrix correction (Bammer et al., 2003)
@@ 2.8.2 Experiment 2: Rician noise bias
The effect of Rician noise bias was investigated by introducing noise sampled from a Rician distribution (SNR = [20,50]), rather than Gaussian. As in experiment 1, only $l_{max}=2$ was considered. Parameter estimation was performed using both a standard MSE loss and a Rician likelihood loss (Parker et al., 2025), and the resulting estimations were compared to the ground truth. The integral in the SM forward model is calculated with the analytical approach.
@@ 2.8.3 Experiment 3: Qualitative comparison on in vivo data
To test the INR on in vivo data, SM parameter estimation was executed on the dataset from section 2.7 using the Rician likelihood loss. The MPPCA map for the SMI fitting was estimated on the unprocessed data and b-value up to 1300 s/mm2. Only $l_{max}=2$ was considered. As the acquisition protocol contained negative $b_{\Delta}$ values, the SM forward model was calculated following the numerical integration approach. The results are compared to the estimates from the methods in section 2.5.
@@ 2.8.4 Experiment 4: Estimation of SH order up to $l_{max}=8$
The performance of the proposed method to model higher order FODs is investigated by fitting the INR with SH orders of $l_{max}=\{2,4,6,8\}$, for four different datasets. The noiseless synthetic dataset and the SNR 50 and 20 synthetic data provide insight in the accuracy of FOD estimation in noisy data, while the in vivo dataset qualitatively shows the capability of the INR to estimate higher order FODs on realistic datasets.
@@ 2.8.5 Experiment 5: Effect of gradient non-uniformity correction of SM parameter estimation
The impact of gradient non-uniformities on SM parameter estimation was assessed using in vivo data. For each voxel, the b-value, $b_{\Delta}$, and b-vectors were recalculated to account for scanner-specific gradient deviations following (Bammer et al., 2003). These corrected effective acquisition parameters were then used to fit the model. This analysis was performed for $l_{max}=2$. The differences between the corrected and uncorrected parameter estimates were subsequently evaluated. Gradient non-uniformity correction was implemented using MSE loss function and compared to gradient non-uniformity correction with NLLS.
@@ 2.8.6 Experiment 6: Implicit neural representation for spatial interpolation
To obtain more detailed insight into the continuous spatial representation of the dataset provided by an INR, the INR fit on the Gaussian noise, SNR 50 synthetic dataset at $l_{max}=2$ is sampled at 8x the original resolution in every dimension, resulting in 0.25mm isotropic voxels. The parameter map of $p_{2}$ was visualized in the coronal and sagittal plane for the original resolution output of the model and the linear, cubic and INR upsampling.
@@ 3.1 Quantitative comparison on simulated data
Parameter estimation on simulated data without noise and with SNR = 50 can be found in the supplementary information, showing a less prominent, but still evident improvement in $\rho$ and RMSE of the INR method compared to other estimation methods for all parameters. Fitting without noise shows that the ground truth generating process is not positively biased towards parameter estimates of the INR.
@@ 3.2 Rician noise bias
The correlation plots between the INR with MSE or Rician loss and the ground truth parameters, alongside the parameter maps for both approaches can be seen in Figure 4. For MSE, the Rician bias is visible in the scatterplots and parameter maps as a general over- or underestimation across all voxels. The Rician loss is able to correct the bias, resulting in better correlation with the ground truth. The bias is most significant for $D_{i}$ and $D_{e}$. The bias effect is less prominent when considering SNR = 50, see supplementary information Figure 4.
@@ 3.4 Estimation of SH order up to $l_{max}=8$
The INR outputs for the different SH order FODs of the synthetic datasets are visualized in Figure 6. Qualitative inspection of the FODs shows plausible FOD shapes and directions throughout all datasets and SH orders. Furthermore, there are no notable differences between the FOD estimates for the noiseless and SNR 50 synthetic dataset, indicating noise robustness in the estimate. Small spurious peaks appear in the SNR 20 synthetic dataset, but the fiber orientations indicated by the larger peaks remains almost identical to both comparisons. In the in-vivo dataset the INR produces plausible FOD shapes and directions as well, as visualized in Figure 7. The backgrounds in Figure 6 and 7 show no large voxel-wise differences for $f_{i}$ across the different orders. This holds true for all kernel parameter estimates, which is shown in further detail in the supplementary information section 6.
Figure 6: Visualization of the centrum semi-ovale for different SH orders $l_{max}$ (rows) and datasets (columns), with parameter map $f_{i}$ as background. FODs are scaled for visibility.
Figure 7: Visualization of the centrum semi-ovale for different SH orders $l_{max}$ of the in-vivo dataset, with parameter map $f_{i}$ as background. FODs are scaled for visibility.
@@ 3.5 Effect of gradient non-uniformity correction of SM parameter estimation
Brain parameter maps with and without gradient non-uniformity correction on in vivo data are presented in Fig. 8. Difference maps show a significant effect on $D_{i}$, $D_{e}$. Corrected maps show effect at the edges of the brain, mainly in the frontal lobe. This is to be expected as gradient non-uniformity is strongest there. $D_{i}$ and $D_{e}$ show lower values after correction. The influence of the correction is least apparent on $f_{i}$ and $p_{2}$. The effect of the gradient non-uniformity correction is similar for both INR and NLLS fitting. Lower diffusivity values at the front and back of the brain appear using both methods, as well as higher $p_{2}$ values. The parameter $f$ shows only small differences between the approaches: NLLS shows no effect, while INR shows small corrections throughout the brain. Results of combining Rician bias loss with gradient non-uniformity correction can be found in the supplementary information figure 6.
Figure 8: Effect of gradient non-uniformity correction on the estimation of SM parameters. Top row shows parameter maps without correction. Middle rows show the effect of gradient non-uniformity correction using the INR method. Bottom rows show effect of gradient non-uniformity correction on NLLS parameter estimation. The difference maps are computed relative to the parameter estimates obtained with the same method, but without applying gradient non-uniformity correction.
@@ 4.1 Implications of the results
In this work, we show how INRs can be used to estimate continuous, noise-robust SM parameter maps of simulated and in vivo datasets and with different SH orders. For high SNR levels, the supervised NN method achieves performance metrics close to those of the INR approach; however, it can be significantly more time-consuming due to its reliance on NLLS for generating part of the training data and the need to retrain the model for each specific acquisition protocol (see supplementary Fig. 3). At higher noise levels (SNR = 20), the INR method clearly outperforms all other methods (see Fig. 3). On in vivo data the underlying representation shows a more structurally correlated appearance, without large inter-voxel variability. This is further illustrated by upsampling the INR at high resolution. Parameter estimates for FOD up to SH order eight can be provided alongside the other SM parameters without introducing bias in other parameters as shown in the supplementary information. The self-supervised (Lim et al., 2022; Sen et al., 2024) nature of the method avoids training-set bias prevalent in supervised fitting methods.
@@
The proposed hyperparameters $n_{p}=5000$ and $n_{h}=2048$ provide a robust setting that can provide good representations of dMRI datasets with different sizes, acquisition protocols, and levels of noise. An exploration of these hyperparameters can be found in the supplement of (Hendriks et al., 2025). The hyperparameter $\sigma^{2}$ provides a convenient way of tuning the model to provide stronger or weaker spatial regularization, as detailed in the supplementary information section 1 which shows results for $\sigma^{2}=1$ (extremely smooth) up to $\sigma^{2}=8$ (granular).
@@
Fitting time is a critical factor when applying dMRI microstructure modeling, and various efforts have been made to speed up the computationally heavy nonlinear optimization and enable large-scale population studies (Daducci et al., 2015a; Harms et al., 2017; Diao and Jelescu, 2023). INRs circumvent this through its inherent self-supervised MLP structure, which allows for efficient, continuous representation of the parameter space without requiring voxel-wise optimization. The INRs are fit on consumer grade hardware in around 5 up to at most 23 minutes for the scenarios tested, much faster than classic NLLS approaches. Using the analytical integration approach decreases training time considerably, possible when excluding negative $b_{\Delta}$ values in acquisition protocols. Once the INR is fit to the dataset, the inference time is negligible. For example, an $lmax=2$ model with $n_{h}=2048$ and $n_{p}=5000$ can perform inference at one million coordinates in 2.7 seconds, including data writing times to and from the GPU. However, since INRs require a model to be fit to every individual subject, a supervised learning approach could remain faster for datasets consisting of many subjects with identical acquisition protocols, despite the considerable amount of training time it requires initially (e.g. 83 minutes on GPU excluding initial NLLS parameter estimations, for the supervised NN method (de Almeida Martins et al., 2021)). Nonetheless, the duration of the fitting process for self-supervised learning in combination with dMRI models has been considerably shortened when applying transfer learning (Li et al., 2024) or hash-encodings (Dwedari et al., 2024), which could make on-the-fly fitting of INRs possible.
@@
The method’s ability to incorporate gradient non-uniformity correction in the fitting process provides an advantage over typical supervised methods, for which the training set would be impractically large to capture the spatial variability. This correction is essential as even small non-uniformities can affect parameter maps (Bammer et al., 2003; Mesri et al., 2020; Mohammadi et al., 2012; Morez et al., 2021), and the availability of high-performance gradient coils suffering from significant gradient non-uniformities is increasing (Jones et al., 2018). To our knowledge, SMI is the only framework that has incorporated gradient non-uniformity correction into the SM fitting process, in the form of PIPE (Coelho et al., 2025). This approach uses SVD and linear regression to approximate the exact acquisition settings in each voxel for linear tensor encoding (LTE) acquisitions.
4.2 Limitations of the work
A limitation of using simulated data for evaluation is the variety of possible approaches to generating the ground truth. In this work, the intention was to create a ground truth with structurally smooth characteristics assumed to mimic real brain tissue. However, factors such as voxel size play a role and need to be further investigated. The ground truth generated for the synthetic experiments makes use of SMI for generating the underlying parameter maps. This could potentially bias the parameters to be in a range that favors estimation using SMI. We indeed observed that estimation on the noiseless signal showed optimal performance for SMI and NLLS (see supplementary Fig. 2). The lower performance of the INR method on simulated data without noise arises from its inability to capture the input data perfectly, as it does not model individual voxels, unlike the other methods. Nevertheless, we have attempted to reduce biases that would benefit a particular method by smoothing the parameter maps and using FODs from a different source (MSMT-CSD). Omitting the smoothing still resulted in the highest performance for INR, see supplementary information figure 7. Another limitation related to the ground truth is that the MGH dataset used in this study contained only LTE acquisitions, which may have led to inaccurate parameter estimates (see discussion at the end of this section). We have investigated the impact of other possible sources of severe bias on creating ground truth parameter maps from the MGH dataset, such as the relatively short diffusion time, noise estimation procedure, and included b-values (results not shown). We found similar overall distributions and linear voxel-wise correlations when using longer diffusion times, noise estimate from repeated b=0 s/mm2 images, and excluding b=200 s/mm2 and b¿10.000 s/mm2 images. Nevertheless, creating a ground truth that balances capturing anatomical reality while exerting sufficient control remains an important avenue to further explore.